OpenStudy (anonymous):

Use the arc length formula and the given information to find r. s = 12 cm, θ = 36°; r = ?

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

s = (2 pi r) (36/360) = 12 solve for r

OpenStudy (anonymous):

120/pi? @douglaswinslowcooper

OpenStudy (anonymous):

60/pi*

OpenStudy (anonymous):

r = (12)(10)/(2 pi) = 60/pi right